Star Videos, Inc., produces short musical videos for sale to retail outlets. The company’s balance sheet accounts as of January 1 are given below.

Star Videos, Inc.
Balance Sheet
January 1
Assets
Cash $ 81,800
Accounts receivable 104,400
Inventories:
Raw materials (film, costumes) $ 17,200
Videos in process 57,400
Finished videos awaiting sale 85,800 160,400
Prepaid insurance 12,650
Studio and equipment (net) 588,000
Total assets $ 947,250
Liabilities and Stockholders’ Equity
Accounts payable $ 203,000
Retained earnings 744,250
Total liabilities and stockholders’ equity $ 947,250

Because the videos differ in length and in complexity of production, the company uses a job-order costing system to determine the cost of each video produced. Studio (manufacturing) overhead is charged to videos on the basis of camera-hours of activity. The company’s predetermined overhead rate for the year ($40 per camera-hour) is based on a cost formula that estimated $280,000 in manufacturing overhead for an estimated allocation base of 7,000 camera-hours. Any underapplied or overapplied overhead is closed to cost of goods sold. The following transactions were recorded for the year:

  1. Film, costumes, and similar raw materials purchased on account, $195,500.
  2. Film, costumes, and other raw materials issued to production, $209,500 (85% of this material was considered direct to the videos in production, and the other 15% was considered indirect).
  3. Utility costs incurred (on account) in the production studio, $85,800.
  4. Depreciation recorded on the studio, cameras, and other equipment, $106,000. Three-fourths of this depreciation related to actual production of the videos, and the remainder related to equipment used in marketing and administration.
  5. Advertising expense incurred (on account), $175,000.
  6. Salaries and wages paid in cash as follows:
Direct labor (actors and directors) $ 89,200
Indirect labor (carpenters to build sets, costume designers, and so forth) $ 82,500
Administrative salaries $ 112,800
  1. Prepaid insurance expired during the year, $9,950 (70% related to production of videos, and 30% related to marketing and administrative activities).
  2. Miscellaneous marketing and administrative expenses incurred (on account), $10,350.
  3. Studio (manufacturing) overhead was applied to videos in production. The company recorded 7,250 camera-hours of activity during the year.
  4. Videos that cost $566,000 to produce according to their job cost sheets were transferred to the finished videos warehouse to await sale and shipment.
  5. Sales for the year totaled $1,000,000 and were all on account.
  6. The total cost to produce the videos that were sold according to their job cost sheets was $611,070.
  7. Collections from customers during the year totaled $950,000.
  8. Payments to suppliers on account during the year, $556,000.
  9. Underapplied or overapplied overhead $__?__.

Required:

1. Prepare a transaction analysis that records all of the above transactions.

2. Prepare a schedule of cost of goods manufactured for the year.

3. Prepare a schedule of cost of goods sold for the year.

4. Prepare an income statement for the year.
Answer #1

1.

Star Videos, Inc.
Transaction Analysis
For the Year Ended December 31
Cash Accounts Receivable Raw Materials Videos in Process Finished Videos Manufacturing Overhead Prepaid Insurance Studio & Equipment (net) = Accounts Payable Retained Earnings
Beginning balance @ 1/1 81800 104400 17200 57400 85800 0 12650 588000 = 203000 744250
a) Raw material purchases 195500 = 195500
b) Raw materials used -209500 178075 31425 =
c) Utility costs 85800 = 85800
d) Depreciation charges 79500 -106000 = -26500
e) Advertising = 175000 -175000
f) Salaries & wages -284500 89200 82500 = -112800
g) Prepaid insurance 6965 -9950 = -2985
h) Miscellaneous marketing = 10350 -10350
i) Applied overhead 290000 -290000 =
j) Transfer completed videos to finished goods -566000 566000 =
k) Sales 1000000 = 1000000
l) Transfer finished goods to cost of goods sold -611070 = -611070
m) Cash collections from customers 950000 -950000 =
n) Payment to suppliers -556000 = -556000
o) Overapplied overhead 3810 = 3810
Ending balances @12/31 191300 154400 3200 48675 40730 0 2700 482000 = 113650 809355

2.

Star Videos, Inc.
Schedule of Cost of Goods Manufactured
For the Year Ended December 31
Beginning raw materials inventory $ 17200
Add: Raw material purchases 195500
Raw materials available for use 212700
Less: Ending raw materials inventory 3200
Raw materials used 209500
Less: Indirect materials 31425 178075
Direct labor 89200
Manufacturing overhead 290000
Total manufacturing costs 557275
Add: Beginning videos in process inventory 57400
Total videos in process inventory 614675
Less: Ending videos in process inventory 48675
Cost of goods manufactured 566000

3.

Star Videos, Inc.
Schedule of Cost of Goods Sold
For the Year Ended December 31
Cost of goods manufactured 566000
Add: Beginning finished goods inventory 85800
Cost of goods available for sale 651800
Less: Ending finished goods inventory 40730
Cost of goods sold 611070
Less: Overapplied overhead 3810
Adjusted cost of goods sold $ 607260

4.

Star Videos, Inc.
Income Statement
For the Year Ended December 31
Sales revenue 1000000
Cost of goods sold 607260
Gross profit 392740
Operating expenses 327635
Net income $ 65105